# Notions of entropy for unitary representations

### v2

@@ 4.3 Functionals on algebras of matrices
The $k^{\mathrm{th}}$ root appears on the left-hand side of (4.10) because we define $\Delta\varphi$ in (4.8) using the unnormalized tracial functional $\tau\otimes\mathrm{tr}_{k}$, but the application of Proposition 3.9 requires the normalized functional $\tau\otimes\overline{\mathrm{tr}}_{k}$. On the right-hand side of (4.10), the corresponding $k^{\mathrm{th}}$ power makes no difference to the condition that $\Delta_{\tau\otimes\mathrm{tr}_{k}}(a)\geq 1$.
@@ 5 Group algebras and positive definite functions
This section specializes some results from the previous two sections to the case when $\mathfrak{A}$ is a group C*-alebra $C^{\ast}\Gamma$ for a countable discrete group $\Gamma$. We write $e$ for the identity in $\Gamma$.
We write $\mathbf{C}[\Gamma]$ for the complex group algebra of $\Gamma$, and regard it concretely as the space of finitely supported functions from $\Gamma$ to $\mathbf{C}$. It is an algebra over $\mathbf{C}$ with convolution as multiplication, and it becomes a $\ast$-algebra under its standard involution: see [Fol16, Section 2.5] or [Dix77, Chapter 13] (which both allow the greater generality of locally compact groups). Given $g\in\Gamma$, we write $\delta_{g}$ for its canonical image in $\mathbf{C}[\Gamma]$, which is the indicator function of the singleton $\{g\}$. The unit of $\mathbf{C}[\Gamma]$ is $\delta_{e}$.
More generally, we may identity $\mathrm{M}_{k,\ell}(\mathbf{C}[\Gamma])$ with the vector space $\mathrm{M}_{k,\ell}[\Gamma]$ of finitely supported maps from $\Gamma$ to $\mathrm{M}_{k,\ell}$. Given $\varphi:\Gamma\to\mathrm{M}_{k,\ell}$ and $\psi:\Gamma\to\mathrm{M}_{\ell,m}$, at least one of them finitely supported, we extend the usual definition of convolution by writing
@@
The resulting map $\varphi\ast\psi$ takes values in $\mathrm{M}_{k,m}$. The individual summands here are matrix products, so the order of factors in each summand matters unless $k=\ell=m=1$. If both $\varphi$ and $\psi$ are finitely supported then so is $\varphi\ast\psi$. If $k=\ell=m$ then this defines the structure of $\mathrm{M}_{k}[\Gamma]$ as a group algebra with matrix coefficients. Given three or more functions on $\Gamma$, all but one of them finitely supported, we can multiply them by repeating (5.1). The resulting product is associative, by the usual proof for the scalar-valued case.
@@
(see [Fol16, Proposition 2.40(a)] for the scalar-valued case).
The group C*-algebra $C^{\ast}\Gamma$ is the maximal C*-completion of $\mathbf{C}[\Gamma]$. Representations of $C^{\ast}\Gamma$ are in one-to-one correspondence with unitary representations of $\Gamma$ itself, and we generally use the same notation for a representation of $C^{\ast}\Gamma$ and for its restriction to $\Gamma$. See [Fol16, Section 7.1] or [Dix77, Section 13.9] for full accounts of this relationship. For each $g\in\Gamma$, we continue to write $\delta_{g}$ for its image in $C^{\ast}\Gamma$. We can identify $\mathrm{M}_{k}(C^{\ast}\Gamma)$ as the corresponding completion of $\mathrm{M}_{k}[\Gamma]$ in the same way.
@@
In representation theory, the function on $\Gamma$ given by $\langle\pi(\cdot)v,u\rangle$ is called the $(u,v)$-matrix element of the representation.
@@
In the most common case $k=1$, the scalar-valued functions on $\Gamma$ that arise this way are also sometimes called ‘functions of positive type’. This case appears first in work of Gelfand and Raikov [GR43], and can also be found in [Fol16, Section 3.3] or [Dix77, Section 13.4]. It is an analog for groups of applying the GNS construction to a state on a C*-algebra. Similarly, the generalization to $\mathrm{M}_{k}$-valued positive definite functions on a group is a cousin of Stinespring’s theorem. For matrix-valued positive definite functions on groups, the essential idea of the construction appears in [Neu43], although it is presented there only for Abelian groups. It is fully covered by the generality in [Kun67], as are various further extensions.
So there is a bijection between the set of $\mathrm{M}_{k}$-valued positive definite maps on $\Gamma$ and the space $\mathfrak{B}(C^{\ast}\Gamma,\mathrm{M}_{k})_{+}$. Under this bijection, the weak topology on $\mathfrak{B}(C^{\ast}\Gamma,\mathrm{M}_{k})_{+}$ corresponds to the usual weak topology of $\ell^{\infty}(\Gamma;\mathrm{M}_{k})$, and when restricted to uniformly bounded subsets it coincides with the topology of pointwise convergence.
Let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite, and let $a\in\mathrm{M}_{k}[\Gamma]$ regarded as an element of $\mathrm{M}_{k}(C^{\ast}\Gamma)$. Having defined convolution in the generality of (5.1), we may use it to express the pairing from (4.3) like this:
$\langle\varphi,a\rangle=\sum_{g}\frac{1}{k}\sum_{ij}a_{ij}(g)\varphi_{ij}(g)=\sum_{g}\overline{\mathrm{tr}}_{k}(\widehat{\varphi}(g^{-1})a(g))=\overline{\mathrm{tr}}_{k}((\widehat{\varphi}\ast a)(e)),$ (5.5)
@@
$\widehat{\varphi}(g):=\varphi(g^{-1})^{\mathrm{T}}\qquad(g\in\Gamma).$ (5.6)
That is, we apply inversion in $\Gamma$ and transpose to elements of $\mathrm{M}_{k}$. These conventions loosely generalize the conjugate that appears in the usual definition of the Fourier–Stieltjes transform of a finite measure on a locally compact Abelian group [Fol16, Section 4.2].
@@
More advanced accounts of unitary group representations and their relation to the group C*-algebras can be found in [BdlHV08, Appendices A–C][BdlH20, Chapters 1 and 2], or [FD88, Chapter VI].
The construction of a representation from a matrix-valued positive definite function on a group can be found as [Pau02, Theorem 4.8]. The relationship between positive definite functions on a group and completely positive maps on the group C*-algebra is discussed further at the end of [Pau02, Chapter 4].
@@ 6 A Szegő-like theorem over amenable groups
This section proves Theorem A. The proof is split across Subsections 6.1 and 6.2. Subsection 6.3 offers some comparisons with previous work and other discussion.
@@ 6.1 Lower bound
For example, if $k=1$, then $x_{1}$ is a tracial vector for $\widetilde{\tau}$, and the desired inequality is simply (3.11) for $T^{2}$. Put roughly, we prove the general case by choosing carefully a single vector in the $k$-fold tensor product $H^{\otimes k}$ that reduces the desired inequality to this special case.
@@
where $A$ is in the $i^{\mathrm{th}}$ position. The operators $A_{1}$, …, $A_{k}$ are all still bounded, they commute, and if $A$ is self-adjoint then so is every $A_{i}$. For another bounded operator $T$ on $H$, the tensor product $T^{\otimes k}\in\mathfrak{B}(H^{\otimes k})$ is equal to the product $T_{1}T_{2}\cdots T_{k}$.
@@
$x_{1}\wedge\cdots\wedge x_{k}:=\frac{1}{\sqrt{k!}}\sum_{\pi\in S_{k}}\mathrm{sgn}(\pi)x_{\pi(1)}\otimes\cdots\otimes x_{\pi(k)}\qquad(x_{1},\dots,x_{k}\in H),$ (6.1)
where $S_{k}$ is the symmetric group on $\{1,2,\dots,k\}$. This vector is called the alternating product of $x_{1}$, …, $x_{k}$. Since $H$ is a Hilbert space, the closed span of all alternating product vectors can be identified with the alternating product space $H^{\wedge k}$: see, for instance, [Sim05c, Section 1.5]. If $x_{1}$, …, $x_{k}$ are orthonormal then $x_{1}\wedge\cdots\wedge x_{k}$ is a unit vector, and more generally alternating products satisfy
$\langle x_{1}\wedge\cdots\wedge x_{k},y_{1}\wedge\cdots\wedge y_{k}\rangle=\det[\langle x_{i},y_{j}\rangle]:$ (6.2)
see, for instance, [Sim05c, equation (1.10)].
@@
where we use Jensen’s inequality [Rud87, Theorem 3.3] on the second line and (6.2) on the last line. Combining this calculation with (6.4) completes the proof.
@@
Now let $\Gamma$ be any countable group, let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite. As in Theorem A, for any finite subset $F$ of $\Gamma$, we consider the $F$-by-$F$ block matrix
@@
The Lebesgue decomposition from Proposition 4.6 gives $\varphi[F_{n}]\geq_{\mathrm{pd}}\varphi_{\mathrm{ac}}[F_{n}]$ for every $n$, and so their determinants are ordered the same way [HJ13, Corollary 7.7.4(e)]. We may therefore discard $\varphi_{\mathrm{sing}}$ and assume that $\varphi$ is $\lambda$-normal.
Let $\lambda$ be the left regular representation of $\Gamma$, and let $H:=\ell^{2}(\Gamma)$ and $\xi:=\delta_{e}$. Denote by $\widetilde{\tau}$ the resulting normal tracial state on either $\lambda(\Gamma)^{\prime\prime}$ or $\lambda(\Gamma)^{\prime}$ from (3.5). Let $k$ be a positive integer, let $\xi_{1}$, …, $\xi_{k}$ be the cyclic $k$-tuple for $\lambda^{\oplus k}$ as in (4.7), and let $\mathfrak{M}:=\lambda^{\oplus k}(\Gamma)^{\prime}$, which we identify with $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime})$. Then this $k$-tuple satisfies
@@
Regarded as a unitary embedding from $\mathbf{C}^{\oplus k|F|}$ into $H$, this satisfies

### v3

@@ 4.3 Functionals on algebras of matrices
The $k^{\mathrm{th}}$ root appears on the left-hand side of (4.13) because we define $\Delta\varphi$ in (4.11) using the un-normalized tracial functional $\tau\otimes\mathrm{tr}_{k}$, but the application of Proposition 3.9 during the proof requires the normalized functional $\tau\otimes\overline{\mathrm{tr}}_{k}$. On the right-hand side of (4.13), the corresponding $k^{\mathrm{th}}$ power makes no difference to the condition that $\Delta_{\tau\otimes\mathrm{tr}_{k}}(a)\geq 1$.
For matrix-valued completely positive maps, Proposition 4.13 shortens the proof of the following.
Corollary 4.14.
Let $\lambda$ and $\tau$ be as above, let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$, and let $\theta$ be any joining of $\varphi$ and $\psi$ (recall Definition 4.4). Then
$\Delta\theta_{\mathrm{ac}}\leq\Delta\varphi_{\mathrm{ac}}\cdot\Delta\psi_{\mathrm{ac}}.$
Equality holds if $\theta_{\mathrm{ac}}=\mathrm{diag}(\varphi_{\mathrm{ac}},\psi_{\mathrm{ac}})$ (and so, in particular, if $\theta=\mathrm{diag}(\varphi,\psi)$).
Proof.
Step 1. Suppose that $T_{1}$ and $T_{2}$ are non-negative square-integrable operators affiliated to $\lambda^{\oplus k}(\mathfrak{A})^{\prime}$ and $\lambda^{\oplus\ell}(\mathfrak{A})^{\prime}$, respectively. Let $T:=T_{1}\oplus T_{2}$, so this is a non-negative square-integrable operator affiliated to $\lambda^{\oplus(k+\ell)}(\mathfrak{A})^{\prime}$. If $E_{1}$ and $E_{2}$ are the respective spectral resolutions of $T_{1}$ and $T_{2}$, then the spectral resolution of $T$ is given by
$E(A):=E_{1}(A)\oplus E_{2}(A)\qquad\hbox{for any Borel}\ A\subset[0,\infty).$
This gives the calculation of the Fuglede–Kadison determinant of $T$:
$\log\Delta_{\widetilde{\tau}\otimes\mathrm{tr}_{k+\ell}}T$ $=\int_{[0,\infty)}\log t\ (\widetilde{\tau}\otimes\mathrm{tr}_{k+\ell})(E(dt))$
$=\int_{[0,\infty)}\log t\ \big{(}(\widetilde{\tau}\otimes\mathrm{tr}_{k})(E_{1}(dt))+(\widetilde{\tau}\otimes\mathrm{tr}_{\ell})(E_{2}(dt))\big{)}$
$=\log\Delta_{\widetilde{\tau}\otimes\mathrm{tr}_{k}}T_{1}+\log\Delta_{\widetilde{\tau}\otimes\mathrm{tr}_{\ell}}T_{2}.$
Exponentiating, and then applying this identity to the operators that represent $\varphi_{\mathrm{ac}}$ and $\psi_{\mathrm{ac}}$ according to Proposition 4.10, we arrive at the desired equality when $\theta_{\mathrm{ac}}$ equals $\mathrm{diag}(\varphi_{\mathrm{ac}},\psi_{\mathrm{ac}})$.
Step 2. Now consider an arbitrary joining $\theta$. Pick any positive and invertible elements $a_{1}\in\mathrm{M}_{k}(\mathfrak{A})$ and $a_{2}\in\mathrm{M}_{\ell}(\mathfrak{A})$ that satisfy $\Delta_{\tau\otimes\mathrm{tr}_{k}}(a_{1})\geq 1$ and $\Delta_{\tau\otimes\mathrm{tr}_{\ell}}(a_{2})\geq 1$. Pick also a positive real value $r$, and let $a:=\mathrm{diag}(r^{\ell}a_{1},r^{-k}a_{2})$, so this lies in $\mathrm{M}_{k+\ell}(\mathfrak{A})$. By the calculation in Step 1, we have
$\Delta_{\tau\otimes\mathrm{tr}_{k+\ell}}(a)=\Delta_{\tau\otimes\mathrm{tr}_{k}}(r^{\ell}a_{1})\cdot\Delta_{\tau\otimes\mathrm{tr}_{\ell}}(r^{-k}a_{2})=\frac{r^{k\ell}}{r^{k\ell}}\Delta_{\tau\otimes\mathrm{tr}_{k}}(a_{1})\cdot\Delta_{\tau\otimes\mathrm{tr}_{\ell}}(a_{2})\geq 1.$
On the other hand,
$\langle\theta,a\rangle=\frac{kr^{\ell}}{k+\ell}\langle\varphi,a_{1}\rangle+\frac{\ell r^{-k}}{k+\ell}\langle\psi,a_{2}\rangle.$
Taking the infimum over $a_{1}$ and $a_{2}$, this expression becomes
$\frac{kr^{\ell}}{k+\ell}(\Delta\varphi_{\mathrm{ac}})^{1/k}+\frac{\ell r^{-k}}{k+\ell}(\Delta\psi_{\mathrm{ac}})^{1/\ell}$
by Proposition 4.13. Finally, by the inequality of arithmetic and geometric means, the infimum of this last expression over $r$ is equal to $(\Delta\varphi_{\mathrm{ac}}\cdot\Delta\psi_{\mathrm{ac}})^{1/(k+\ell)}$. By Proposition 4.13 again, this gives the desired upper bound on $\Delta\theta_{\mathrm{ac}}$. ∎
@@
The textbook [117] is dedicated to completely positive and completely bounded maps between operator algebras, and includes a more complete guide to original references. The basic results that we need are mostly covered in Chapters 4 and 6 of that book. Some more recent uses of completely positive maps in the study of C-algebras, including group C-algebras, can be found in [30, Sections 1.2 and 2.5 and Appendix D].
@@ 5 Group algebras and positive definite functions
This section specializes some results from the previous two sections to the case when $\mathfrak{A}=C^{\ast}\Gamma$ for a countable discrete group $\Gamma$. We write $e$ for the identity in $\Gamma$.
We write $\mathbf{C}[\Gamma]$ for the complex group algebra of $\Gamma$, and regard it concretely as the space of finitely supported functions from $\Gamma$ to $\mathbf{C}$. It is an algebra over $\mathbf{C}$ with convolution as multiplication, and it becomes a $\ast$-algebra under its standard involution: see [62, Section 2.5] or [48, Chapter 13] (which both allow the greater generality of locally compact groups). Given $g\in\Gamma$, we write $\delta_{g}$ for its canonical image in $\mathbf{C}[\Gamma]$, which is the indicator function of the singleton $\{g\}$. The unit of $\mathbf{C}[\Gamma]$ is $\delta_{e}$.
More generally, we may identity $\mathrm{M}_{k}(\mathbf{C}[\Gamma])$ with the vector space $\mathrm{M}_{k}[\Gamma]$ of finitely supported maps from $\Gamma$ to $\mathrm{M}_{k}$. Given $\varphi,\psi:\Gamma\to\mathrm{M}_{k}$, at least one of them finitely supported, we extend the usual definition of convolution by writing
@@
The resulting map $\varphi\ast\psi$ takes values in $\mathrm{M}_{k}$. The individual summands here are matrix products, so the order of factors in each summand matters unless $k=1$. If both $\varphi$ and $\psi$ are finitely supported then so is $\varphi\ast\psi$, and then (5.1) defines the structure of $\mathrm{M}_{k}[\Gamma]$ as a group algebra with matrix coefficients. Given three or more functions on $\Gamma$, all but one of them finitely supported, we can multiply them by repeating (5.1). The resulting product is associative, by the usual proof for the scalar-valued case.
@@
(see [62, Proposition 2.40(a)] for the scalar-valued case).
The group C*-algebra $C^{\ast}\Gamma$ is the maximal C*-completion of $\mathbf{C}[\Gamma]$. Representations of $C^{\ast}\Gamma$ are in one-to-one correspondence with unitary representations of $\Gamma$ itself, and we generally use the same notation for a representation of $C^{\ast}\Gamma$ and for its restriction to $\Gamma$. See [62, Section 7.1] or [48, Section 13.9] for full accounts of this relationship. For each $g\in\Gamma$, we continue to write $\delta_{g}$ for its image in $C^{\ast}\Gamma$. We can identify $\mathrm{M}_{k}(C^{\ast}\Gamma)$ as the corresponding completion of $\mathrm{M}_{k}[\Gamma]$ in the same way.
@@
In representation theory, the function on $\Gamma$ given by $\langle\pi(\cdot)v,u\rangle$ is called the $(u,v)$-matrix element of the representation $\pi$.
@@
In case $k=1$, the scalar-valued functions on $\Gamma$ that arise this way are also sometimes called ‘functions of positive type’. This case appears first in work of Gelfand and Raikov [69], and can also be found in [62, Section 3.3] or [48, Section 13.4]. It is an analog for groups of applying the GNS construction to a state on a C*-algebra. Similarly, the generalization to $\mathrm{M}_{k}$-valued positive definite functions on a group is a cousin of Stinespring’s theorem. For matrix-valued positive definite functions on groups, the essential idea of the construction appears in [116], although it is presented there only for Abelian groups. It is fully covered by the generality in [99], as are various further extensions.
So there is a bijection between the set of $\mathrm{M}_{k}$-valued positive definite maps on $\Gamma$ and the space $\mathfrak{B}(C^{\ast}\Gamma,\mathrm{M}_{k})_{+}$. Under this bijection, the weak topology on $\mathfrak{B}(C^{\ast}\Gamma,\mathrm{M}_{k})_{+}$ corresponds to the usual weak topology restricted from $\ell^{\infty}(\Gamma;\mathrm{M}_{k})$, and when restricted further to any uniformly bounded subset it coincides with the topology of pointwise convergence.
Let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite, let $a\in\mathrm{M}_{k}[\Gamma]$, and regard $a$ as an element of $\mathrm{M}_{k}(C^{\ast}\Gamma)$. Having defined convolution in the generality of (5.1), we may use it to express the pairing from (4.7) like this:
$\langle\varphi,a\rangle=\frac{1}{k}\sum_{ij}\sum_{g}a_{ij}(g)\varphi_{ij}(g)=\sum_{g}\overline{\mathrm{tr}}_{k}(\widehat{\varphi}(g^{-1})a(g))=\overline{\mathrm{tr}}_{k}((\widehat{\varphi}\ast a)(e)),$ (5.5)
@@
$\widehat{\varphi}(g):=\varphi(g^{-1})^{\mathrm{T}}\qquad(g\in\Gamma)$ (5.6)
(that is, we apply inversion in $\Gamma$ and transpose to elements of $\mathrm{M}_{k}$). The conventions in (5.6) loosely generalize the conjugate that appears in the usual definition of the Fourier–Stieltjes transform of a finite measure on a locally compact Abelian group [62, Section 4.2].
@@
Thorough accounts of unitary group representations and their relation to group C*-algebras can be found in [15, Appendices A–C][14, Chapters 1 and 2], or [61, Chapter VI].
The construction of a representation from a matrix-valued positive definite function on a group can be found as [117, Theorem 4.8]. The relationship between positive definite functions on a group and completely positive maps on the group C*-algebra is discussed further at the end of [117, Chapter 4].
@@ 6 A Szegő-like theorem over amenable groups
This section proves Theorem A. The proof is split across Subsections 6.1 and 6.2. Subsection 6.3 offers comparisons with previous work and some other discussion.
@@ 6.1 Lower bound
For example, if $k=1$, then $x_{1}$ is a tracial vector for $\widetilde{\tau}$, and the desired inequality is simply (3.11) for $T^{2}$. Put roughly, we prove the general case by choosing carefully a single vector in the tensor product $H^{\otimes k}$ that reduces the desired inequality to this special case.
@@
where $A$ is in the $i^{\mathrm{th}}$ position. The operators $A_{1}$, …, $A_{k}$ are all still bounded; they commute; and if $A$ is self-adjoint then so is every $A_{i}$. For another bounded operator $T$ on $H$, the tensor product $T^{\otimes k}\in\mathfrak{B}(H^{\otimes k})$ is equal to the product $T_{1}T_{2}\cdots T_{k}$.
@@
$x_{1}\wedge\cdots\wedge x_{k}:=\frac{1}{\sqrt{k!}}\sum_{\pi}\mathrm{sgn}(\pi)x_{\pi(1)}\otimes\cdots\otimes x_{\pi(k)}\qquad(x_{1},\dots,x_{k}\in H),$ (6.1)
where the sum runs over all permutations of $\{1,2,\dots,k\}$. This vector is called the alternating product of $x_{1}$, …, $x_{k}$. Since $H$ is a Hilbert space, the closed span of all alternating product vectors can be identified with the alternating product space $H^{\wedge k}$: see, for instance, [138, Section 1.5]. If $x_{1}$, …, $x_{k}$ are orthonormal then $x_{1}\wedge\cdots\wedge x_{k}$ is a unit vector, and more generally alternating products satisfy
$\langle x_{1}\wedge\cdots\wedge x_{k},y_{1}\wedge\cdots\wedge y_{k}\rangle=\det[\langle x_{i},y_{j}\rangle]$ (6.2)
(see, for instance, [138, equation (1.10)]).
@@
where we use Jensen’s inequality [126, Theorem 3.3] on the second line and (6.2) on the last line. Combining this calculation with (6.4) completes the proof.
@@
Now let $\Gamma$ be any countable group and let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite. As in Theorem A, for any finite subset $F$ of $\Gamma$, we consider the $F$-by-$F$ block matrix
@@
The Lebesgue decomposition from Proposition 4.9 gives $\varphi[F_{n}]\geq\varphi_{\mathrm{ac}}[F_{n}]$ in the positive definite ordering for every $n$, and so their determinants are ordered the same way [88, Corollary 7.7.4(e)]. We may therefore discard $\varphi_{\mathrm{sing}}$ and assume that $\varphi$ is $\lambda$-normal.
Let $\lambda$ be the left regular representation of $\Gamma$, and let $H:=\ell^{2}(\Gamma)$ and $\xi:=\delta_{e}$. Denote by $\widetilde{\tau}$ the resulting normal tracial state from (3.5) on either $\lambda(\Gamma)^{\prime\prime}$ or $\lambda(\Gamma)^{\prime}$. Let $k$ be a positive integer, let $\xi_{1}$, …, $\xi_{k}$ be the cyclic $k$-tuple for $\lambda^{\oplus k}$ as in (4.10), and let $\mathfrak{M}:=\lambda^{\oplus k}(\Gamma)^{\prime}$, which we identify with $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime})$. Then this $k$-tuple satisfies
@@
Regarded as a unitary embedding from $\mathbf{C}^{\oplus k|F|}$ into $H$, this tuple satisfies